How many grams of sodium nitrate will be produced if you start with 135.0 grams of sodium hydroxide?


1
Expert's answer
2021-04-12T00:04:16-0400

HNO3 +NaOH->NaNO3 + H2O

1 mole of NaOH = 1 mole of NaNO3

molar mass of NaOH = 39.997 g/mol

Molar Mass of NaNO3 = 84.9947 g/mol

135 g of NaOH = 135/39.997

= 3.375 moles

= 3.375 × 84.9947 = 286.8786g of NaNO3